OpenStudy (anonymous):

A weather balloon is inflated to a volume of 2.2 x 103 L with 37.4 g of helium. What is the density of helium in grams per liter? A) 59 g/L B) 1.7 x 10-2 g/L C) 44.4 g/L

OpenStudy (anonymous):

\[2.2 \times 10^3 liters = 2200 liters\] \[{ 37.4g \over 2200l } = { 0.017 } = 1.7 \times 10^{-2} g/l\]